# Survey of Load-Altering Attacks Against Power Grids: Attack Impact, Detection and Mitigation

### v1

@@ Abstract
The growing penetration of IoT devices in power grids despite its benefits, raises cyber security concerns. In particular, load-altering attacks (LAAs) targetting high-wattage IoT-controllable load devices pose serious risks to grid stability and disrupt electricity markets. This paper provides a comprehensive review of LAAs, highlighting the threat model, analyzing its impact on transmission and distribution networks, and the electricity market dynamics. We also review the detection and localization schemes for LAAs that employ either model-based or data-driven approaches, with some hybrid methods combining the strengths of both. Additionally, mitigation techniques are examined, focusing on both preventive measures, designed to thwart attack execution, and reactive methods, which aim to optimize responses to ongoing attacks. We look into the application of each study and highlight potential streams for future research in this field.
@@ I Introduction
Demand flexibility (DF) is seen as a key element in achieving decarbonization and net-zero goals set by several nations across the globe. Shifting the load to off-peak periods and aligning them with renewable energy sources (RES) can significantly relieve network congestion, thus avoiding costly infrastructure upgrades. Motivated by these considerations, there has been a rapid proliferation of smart energy appliances, such as smart electric vehicle charging stations (EVCS) and smart heat pumps. The connectivity features can be used to schedule their usage, hence offering DF. The penetration of such appliances is expected to grow significantly in the future. For instance, in the United Kingdom (UK), the government is backing the expansion of public EV charging infrastructure to reach $300k$ public chargers by 2030 with an investment totalling £$1.6$b [1], while the capital of Saudi Arabia, Riyadh, targets towards rising the EV fleet to over $1$m by 2030 [2].
However, these devices, collectively referred to as IoT-enabled smart energy appliances, can become a serious source of security vulnerability against the power grid. Several security issues, such as insecure authentication, lack of firmware security upgrades, etc., have already been discovered in these devices (see Section II and Table I for a detailed discussion). They can be particularly attractive targets for malicious actors due to the following factors: (i) the security standards for these devices are still in their infancy, and there is a lack of a unified approach to this end; (ii) they are typically manufactured and installed by third parties, who may typically ignore security features/recommendations due to cost considerations; (iii) the power grid operator does not have direct control over these devices; (iv) they are operated by end-user customers who typically lack security awareness.
These threats go beyond mere theoretical constructs. Real-world cyberattacks involving smart energy appliances have already been witnessed. For instance, an EVCS in the Isle of Wight region was hacked to display inappropriate images [3]. There have also been reports of compromising and disabling EVCS near the Moscow region and hackers using their displays to show political messages [4]. In the UK, EV chargers from a specific vendor were pulled out of the market due to inadequate cybersecurity features in its design, with the concern that the chargers can be used as a means to target the power grid [5]. In December 2023, a US court-authorized operation thwarted a botnet attack targeting several small office/home office routers against a “campaign to target critical infrastructure organizations in the US” [6].
@@
The focus of this work is large-scale compromises, such as a botnet attack targeting a large number of IoT-load devices. A popular instance is the Mirai botnet attack which at its peak targeted $600k$ IoT devices. A real-world illustration of synchronously turning ON/OFF a large number of consumer electrical appliances can be observed during the so-called “TV Pickup effect” experienced by the UK’s power grid [7]. It refers to large-scale load surges or drops in the load during popular football matches. During half and full-time breaks, it has been observed that customers synchronously turn on electric kettles or open their fridge doors, the aggregate effect of which resulted in several GW of load change, as illustrated in Fig. 1. If such events occur due to a cyber attack, they will remain completely unanticipated, resulting in grid stability risks. Table I represents a list of known common vulnerabilities and exposures (CVE) which provide entry points for adversaries for such attacks and execute LAAs.
@@ I-A Difference From Existing Works
The topic of power grid security has attracted significant interest from the research community in the last decade. A vast majority of the works focus on cyberattacks against the power grid’s supervisory control and data acquisition (SCADA) system. Notable examples include false data injection attacks (FDI), denial of service (DoS) or time delay attacks targeting the data communicated between power grid sensors and the control centers, physical attacks that alter the power grid topology (e.g., remotely opening transmission line circuit breakers), coordinated data and physical attacks, time synchronization attacks against phasor measurement units (PMUs), etc. Several survey papers have been written on this topic [8, 9, 10, 11, 12]. The power grid industry and policymakers have taken note of these threats and developed security solutions to counter them that have already made their way to regulations/standards. For example, the IEC 62351 standard provides security recommendations for smart grid communication protocols such as the IEC 61850 protocol popularly used for power grid SCADA communication (e.g., the requirement to use digital signatures by SHA256 and RSA algorithms to counter FDI or replay-type attacks) [13]. Furthermore, in the research literature, several cyber-physical defense solutions have been proposed, such as strategic protection of sensors to prevent undetectable attacks [14], machine learning (ML)-based approaches to detect cyberattacks [11], and moving target defense to invalidate attacker’s knowledge of the target system and thus making the attacks detectable [15].
@@
In contrast, the focus of this survey paper is LAAs, targeting the end-user smart energy appliances that are installed at the customer end as depicted in Fig. 2. As noted previously, compared to the assets located in the SCADA system, these devices are much easier to compromise from an attacker’s point of view. A few recent survey papers have investigated cyber-threats from end-user devices. For instance, [16] surveys the privacy of smart meters, existing problems, and their solutions. In [17, 18, 19, 20], the main focus is on cyberattacks targetting distributed energy resources (DERs) and their components. The work surveys the security of DERs in both cyber and physical layers and covers the attacks targeting DERs and their communication means while, [18] discusses vulnerabilities and cyber-resiliency of DER-based smart grids. Furthermore, [19] makes a comprehensive overview on attacks on photovoltaic systems and their mitigation methods. The work in [20] provides an extensive survey about existing studies on the cybersecurity of smart inverters of DERs. Overall, these works primarily address threats that affect DERs operations, such as manipulating DER protocols or exploiting inverter-level weaknesses to disrupt energy generation or storage capabilities. Unlike DER-based attacks that target the generation and storage aspects of power systems, LAAs manipulate the load directly through IoT-enabled smart appliances, disrupting the balance between power supply and demand. While DER-based attacks compromise the integrity and control of distributed resources, LAAs exploit vulnerabilities in consumer-side appliances, leading to fluctuations in grid stability and potentially causing severe frequency deviations and cascading failures. This distinction underscores the need for dedicated research on LAAs, as they pose unique threats by targeting end-user devices. With DF via smart load devices playing a key role in decarbonization and net-zero efforts, LAAs are emerging as a critical threat vector. To the best of our knowledge, this work is the first comprehensive review of LAAs against power grids.
It is also important to differentiate LAAs from other attack types, such as load-redistribution attacks (LRAs) [21]. Although both attacks concern the manipulation of load within a power system, the underlying mechanisms and objectives differ. LRAs are instances of FDI attacks where the attacker manipulates the load data measured at transmission buses and reported to the control center via the SCADA system. In contrast, LAAs refer to direct manipulation of the load at end-user devices installed at customer sites, making them instances of control attacks. Unlike LRAs, which aim to disrupt optimal power flow decisions after the system has reached a steady state, LAAs directly impact the balance between generation and demand, potentially causing unsafe frequency fluctuations. To summarize, the contributions of this paper are as follows:
@@
The rest of the paper is organized as follows: Section II discusses the threat modeling, Section III surveys the impact of the attack on grids operation and stability as well as energy markets which is followed by studying attack detection and localization frameworks in Section IV. Furthremore, Section V presents the mitigation methods for LAAs, and finally, Section VI remarks the future research and concludes the paper.
@@ II Threat Modeling
Direct system-end attacks targeting the assets in the SCADA system are the most prevalent form of cyber-threats (e.g., those noted in Section I). Conversely, the deployment of DERs at the consumer level along with the large-scale deployment of high-wattage IoT devices introduces new vulnerabilities, enabling adversaries to orchestrate and execute remote cyberattacks. For instance, DERs leverage IoT technology to synchronize their operational controls with power systems. Notably, devices such as smart inverters, which are consumer electronics driven by firmware-based IoT technology, are particularly prone to cyber-threats. When compromised, these devices can interfere with the functioning of the grid, disrupt the delivery of electricity, and potentially destabilize the entire grid [22, 23, 24].
Unlike traditional power generation facilities, DERs are managed by end-users and aggregators, limiting utility companies’ control over electricity production. The IoT interfaces utilized by these DERs and other high-wattage consumer equipment are particularly vulnerable to cyberattacks due to their complex supply chains and limited resources. Such vulnerabilities make them prime targets for various cyber threats. For instance, insecure remote login credentials can be leveraged to spread malware like the Mirai botnet [25]. Additionally, the process of firmware updating presents another risk for introducing malicious software [26]. Furthermore, Ghost domain name system (DNS) attacks pose a significant threat as they can redirect IoT devices to harmful DNS servers [27]. The security of electricity consumers and DERs is also compromised by the frequent use of default or weak passwords, leading to an increased risk of cyberattacks [28].
This paper examines the threat landscape of indirect consumer-level cyberattacks on the power system via load alternation. Such attacks are possible because DER-based inverters and IoT-based high-wattage devices have communication and control interfaces which do not have as advanced cyber-defense as grid-end systems (e.g., SCADA) and can be exploited with the malicious intent to destabilize the grid. Many attack vectors use a multi-layer threat model by leveraging the complexity and interconnections within the power grid. As shown in previous works [29, 30, 31], attacks can target one or multiple layers of the IoT-connected appliances and DERs. Adversaries may opt to infiltrate any part of the supply chain to inject stealthy trojans (i.e., malicious and deliberate modifications, consisting of a multi-layer model of attack vectors aiming to compromise, control, modify, or disable the system or leak sensitive information) into the IoT firmware/software. The attacks can be either physical (i.e., access required to extract device code or network-based (e.g., remote over-the-air manipulation). LAAs emphasize the cross-layer propagation of such attacks which are able to cause significant load alteration of the system conditions with potentially catastrophic consequences in the operating conditions.
@@ II-A Attacker Model
Using the IoT-related vulnerabilities, an attacker would aim to modify control and operation actions of IoT-load devices under normal conditions to actions that would harm DERs and/or distribution and transmission systems. Following this modification, the production and consumption of each DER/IoT-load are changed, and, as a result, the power demand at those nodes will change. Using this mechanism, the attacker will aim to coordinate actions of compromised devices in the system to maximize potential damage. The attack model of LAAs assumes that the attacker possesses some, but not necessarily exact, knowledge of the system design (to drive the system conditions into activating specifc protection and control actions) (e.g., topology, parameters of lines, buses, generators, etc), which can be obtained from publicly available datasets (e.g., [32, 33, 34, 35]), while further details can be inferred with a high accuracy [36, 37].
@@
Existing literature often formalizes the attack model of LAAs by co-optimizing it with the distribution system operations as given in and adopting a common cybersecurity practice, which is to conservatively assume a strong (omniscient) attacker and a weak (naive) defender (e.g., DSO) [38, 39, 40]. Accordingly, the omniscient attacker has full knowledge of the system, while the naive defender has no detection or defense mechanisms to respond. These assumptions lead to assessing the worst-case impact of the attack. The omniscient attacker can pursue instant and lasting effects on the power system and DERs, given oracular knowledge of the system conditions. The attacker objective can vary including: (i) reducing efficiency of DSO operations by reducing the energy and cost performance of the system, (ii) equipment damage and failures by incurring additional wear-and-tear due to violating operation limits on equipment and intelligently triggering relay protection to cause load shedding and blackouts, (iii) stressing the system during extreme events (e.g., natural disasters or peak conditions) to increase the likelihood of equipment failures.
@@
The model considered in the works above analyzes the worst-case impact of cyberattacks, i.e., under the assumption of the omniscient attacker and naive defender. In practice, it is unrealistic that an attacker can monitor every aspect of DER or power system operations. Therefore, it is realistic to account for the attacker’s limited knowledge of actions of uncompromised DERs and IoT-loads, and the DSO, which prevents the attacker from precisely estimating attack impacts on the grid and from incurring the theoretical maximum damage. The attacker’s limited knowledge (constraints) can be considered in various manners, including:
@@
(i) Statistical modeling of uncompromised DERs/IoTs: using open-source data sets, e.g., Pecan Street [41] or Tracebase data repositories [42], the attacker can construct a statistical model to describe the behavior of uncompromised devices. These repositories have sufficient data to estimate the mean and standard deviation, which can be used to parameterize the uncompromised devices via the Gaussian distribution, or more complex statistical models (e.g., to account for temporal and spatial correlations [43]). Under such statistical models, the threat model of LAAs can be reformulated into different instances of computationally tractable robust optimization models [44], which can be used to formalize and compare different attack objectives under various statistical assumptions, i.e., quantify realistic rather than worst-case attack effects.
@@
(ii) Adversarial Modeling of the Response: building on statistical models, one can consider an adversarial behavior of the DSO, which renders a game-theoretic formulation for DSO-attacker interactions. This adversarial modeling can provide a realistic rather than naive grid response and can be implemented in a static or dynamic manner, where the attacker fully or partially observes the system response and modifies the attack accordingly [45, 46]. This can determine the existence, uniqueness, and properties of the equilibrium for single/multi-stage (sequential) games using computationally scalable algorithmic game theory [47, 48].
@@
(iii) Learning for Partial System Observability: The statistical and adversarial models described above can be supported by considering the limited observability of the system by the attacker (e.g., it can only partially observe power flows, voltages, and other system parameters). In this case, the attacker’s problem can be formulated via reinforcement learning (RL), where the observed nodal parameters are the environmental states and the attack damage is the reward. Using such RL frameworks can lead to continuous learning, instead of the omniscient attacker, in a multi-stage adversarial game.

### v2

@@ Abstract
The growing penetration of IoT devices in power grids despite its benefits, raises cybersecurity concerns. In particular, load-altering attacks (LAAs) targeting high-wattage IoT-controllable load devices pose serious risks to grid stability and disrupt electricity markets. This paper provides a comprehensive review of LAAs, highlighting the threat model, analyzing their impact on transmission and distribution networks, and the electricity market dynamics. We also review the detection and localization schemes for LAAs that employ either model-based or data-driven approaches, with some hybrid methods combining the strengths of both. Additionally, mitigation techniques are examined, focusing on both preventive measures, designed to thwart attack execution, and reactive methods, which aim to optimize responses to ongoing attacks. We look into the application of each study and highlight potential streams for future research.
@@ I Introduction
Power grids are witnessing a rapid proliferation of smart energy appliances, such as smart electric vehicle charging stations (EVCS) and smart heat pumps. The connectivity features of these devices can be used to schedule their usage, hence offering demand flexibility (DF). For instance, in the United Kingdom (UK), the government is backing the expansion of public EV charging infrastructure to reach $300$k public chargers by 2030 with an investment totaling £$1.6$b [1], while the capital of Saudi Arabia, Riyadh, targets towards rising the EV fleet to over $1$m by 2030 [2].
However, these IoT-enabled appliances can become a serious source of security vulnerability to target the power grid. Several security issues, such as insecure authentication, lack of firmware security upgrades, etc., have already been discovered in these devices (see Section II for a detailed discussion). Overall, they can be particularly attractive targets for malicious actors due to the following factors: (i) the security standards for these devices are still in their infancy, and there is a lack of a unified approach for securing them; (ii) they are typically manufactured and installed by third parties, who may typically ignore security features/recommendations due to cost considerations; (iii) the power grid operator does not have direct control over these devices; (iv) they are operated by end-user customers who may lack security awareness.
One major threat emerging from these vulnerabilities is the concept of load-altering attacks (LAAs) [3]. LAAs exploit insecure, high-wattage IoT-enabled devices, such as EVCS, smart thermostats, and industrial loads, to manipulate power consumption on a large scale. By coordinating synchronized load changes, attackers can disrupt the power grid’s frequency stability, induce voltage fluctuations, and, in extreme cases, cause cascading failures or blackouts. Unlike traditional cyberattacks targeting SCADA systems or distributed energy resources (DERs), LAAs manipulate end-user loads directly, which makes them harder to detect and mitigate.
The threats go beyond mere theoretical constructs. Real-world cyberattacks involving smart appliances have already been witnessed. For instance, an EVCS was hacked to display inappropriate images [4]. There have also been reports of compromising and disabling of EVCS near the Moscow region and hackers using their displays to show political messages [5]. In the UK, EV chargers from a specific vendor were pulled out of the market due to inadequate cybersecurity features in its design, with the concern that the chargers can be exploited as a means to target the power grid [6]. In December 2023, a US court-authorized operation thwarted a botnet attack targeting several small office/home office routers against a “campaign to target critical infrastructure organizations in the US” [7].
@@
The focus of this work is large-scale compromises, such as a botnet attack targeting a large number of IoT-load devices i.e., LAAs. A popular instance is the Mirai botnet attack which at its peak targeted $600k$ IoT devices. A real-world illustration of synchronously turning ON/OFF a large number of consumer electrical appliances can be observed during the so-called “TV Pickup effect” experienced by the UK’s power grid [8]. It refers to large-scale load surges or drops in the load during popular football matches. During half-time and full-time breaks, it has been observed that customers synchronously turn on electric kettles or open their fridge doors, the aggregate effect of which resulted in several GW of load change, as illustrated in Fig. 1. If such events occur due to a cyber attack, they will remain unanticipated, resulting in grid stability risks. Table I represents a list of known common vulnerabilities and exposures (CVE) which provide entry points for adversaries to execute such attacks.
@@ I-A Difference From Existing Works
The topic of power grid security has attracted significant interest from the research community in the last decade and several survey papers have been written on this topic [9, 10, 11, 12, 13]. A vast majority of the works focus on cyberattacks against the power grid’s supervisory control and data acquisition (SCADA) system. Notable examples include false data injection attacks (FDI), denial of service (DoS) or time delay attacks targeting the data communicated between power grid sensors and the control centers, physical attacks that alter the power grid topology (e.g., remotely opening transmission line circuit breakers), coordinated data and physical attacks, time synchronization attacks against phasor measurement units (PMUs), etc. The power grid industry and policymakers have taken note of these threats and developed security solutions to counter them and already paved their way to form regulations/standards, such as the IEC 62351 standard which provides security recommendations for the IEC 61850 protocol [14]. Furthermore, in the research literature, several cyber-physical defense solutions are proposed in existing papers, such as strategic protection of sensors to prevent undetectable attacks [15], machine learning (ML)-based approaches to detect cyberattacks [12], and moving target defense to invalidate attacker’s knowledge of the target system and thus making the attacks detectable [16].
@@
While the papers discussed before address the threats against the bulk power grid and centralized generations and their control loops [17, 18, 19]. However, with the high integration of the DER in the grids and the vulnerability of these resources against cyberattacks, recent works have turned their attention to studying attacks on distributed and decentralized generations (similar to LAAs which are attacks on edge devices which are scattered all over the grid). In [20, 21, 22, 23], the main focus is on cyberattacks targeting DERs and their components. Reference [20] surveys the security of DERs in both cyber and physical layers and covers the attacks targeting DERs and their communication means while [21] discusses vulnerabilities and cyber-resiliency of DER-based grids. Furthermore, [22] makes a comprehensive overview on attacks on photovoltaic systems and their mitigation methods. The work in [23] provides an extensive survey of existing studies on the cybersecurity of smart inverters of DERs. Overall, these works primarily address threats that affect DER operations, such as manipulating DER protocols or exploiting inverter-level weaknesses to disrupt generation or storage capabilities. Unlike DER-based attacks that target the generation and storage aspects of power systems, LAAs manipulate the load directly through IoT-enabled smart appliances, disrupting the balance between power supply and demand.
While DER-based attacks compromise the integrity and control of distributed resources, LAAs exploit vulnerabilities in consumer-side appliances, leading to fluctuations in grid stability and potentially causing frequency deviations and cascading failures. This distinction underscores the need for dedicated research on LAAs, as they pose unique threats by targeting end-user devices. To the best of our knowledge, this work is the first comprehensive review of LAAs against power grids.
A few recent survey papers have investigated cyber threats from end-user devices. For instance, [24] surveys the privacy of smart meters. However, according to the lack of a comprehensive review paper specifically focusing on LAAs, this survey paper is dedicated to LAAs, targeting the end-user smart energy appliances that are installed at the customer end as depicted in Fig. 2.
To summarize, the contributions of this paper are as follows:
@@
The rest of the paper is organized as follows: Section II discusses the threat modeling, Section III surveys the impact of the attack followed by studying attack detection and localization frameworks in Section IV. Section V presents the mitigation methods for LAAs and eventually, Section VI remarks on the future research and concludes the paper.
@@ II Threat Modeling
Direct system-end attacks targeting the assets in the SCADA system are the most prevalent form of cyber threats (such as those noted in Section I). Conversely, the deployment of DERs at the consumer level along with the large-scale deployment of high-wattage IoT devices introduces new vulnerabilities, enabling adversaries to orchestrate and execute remote cyberattacks. For instance, DERs leverage IoT technology to synchronize their operational controls with power systems. Notably, devices such as smart inverters are particularly prone to cyber-threats. When compromised, these devices can interfere with the functioning of the grid, disrupt the delivery of electricity, and potentially destabilize the entire grid [25, 26, 27].
Unlike traditional power generation facilities, DERs are managed by end-users and aggregators, limiting utility companies’ control over electricity production. The IoT interfaces utilized by these DERs and other high-wattage consumer equipment are particularly vulnerable to cyberattacks due to their complex supply chains and limited resources. Such vulnerabilities make them prime targets for various cyber threats. For instance, insecure remote login credentials can be leveraged to spread malware like the Mirai botnet [28]. Additionally, the process of updating firmware presents another risk for introducing malicious software [29]. Furthermore, Ghost domain name system (DNS) attacks pose a significant threat as they can redirect IoT devices to harmful DNS servers [30]. The security of electricity consumers and DERs is also compromised by the frequent use of default or weak passwords, leading to an increased risk of cyberattacks.
This paper examines the threat landscape of indirect consumer-level cyberattacks on the power system via load alternation. Such attacks are possible because DER-based inverters and IoT-based high-wattage devices have communication and control interfaces which lack the advanced cyber-defense and can be exploited with the malicious intent to destabilize the grid. Many attack vectors use a multi-layer threat model by leveraging the complexity and interconnections within the power grid. As shown in previous works [31, 32, 33], attacks can target one or multiple layers of the IoT-connected appliances and DERs. Adversaries may opt to infiltrate any part of the supply chain to inject stealthy trojans (i.e., malicious and deliberate modifications, consisting of a multi-layer model of attack vectors aiming to compromise, control, modify, or disable the system or leak sensitive information) into the IoT firmware/software. The attacks can be either physical (i.e., access required to extract device code) or network-based (e.g., remote over-the-air manipulation). LAAs emphasize the cross-layer propagation of such attacks which can cause significant load alteration.
@@ II-A Attacker Model
Using IoT-related vulnerabilities, an attacker would aim to modify control and operation actions of IoT-load devices under normal conditions to actions that would harm DERs and/or distribution and transmission systems. Following this modification, the production and consumption of each DER/IoT-load are changed and, as a result, the power demand at those nodes changes. Using this mechanism, the attacker will aim to coordinate actions of compromised devices in the system to maximize the potential damage. The attack model of LAAs assumes that the attacker possesses some, but not necessarily exact, knowledge of the system design (to drive the system conditions into activating specific protection and control actions) e.g., topology, parameters of lines, buses, generators, etc, which can be obtained from publicly available datasets (e.g., [34, 35, 36, 37]), while further details can be inferred with a high accuracy [38, 39].
First, the knowledge of network topology, line parameters, bus data, generator configurations, and EVCSs can be deduced from publicly available datasets [36]. Examples include collecting locations and power ratings of EVCSs, real-time and historical usage data from EV smartphone applications, locations of generators, transmission lines, transmission substations, and their capacity from public repositories such as the U.S. Energy Information Administration, and real-time generation and demands from NYISO datasets [36]. However, some specifications like the length of lines are estimated, and may not reflect the true values. Using such gathered intelligence, [40] design an (EV-oriented) LAA which constructs the attack vector based on the grid’s frequency only and still causes a destructive impact.
Secondly, topology identification methods allow attackers to deduce network structures based on observable measurements in both transmission and distribution networks [41, 42]. Specifically, in distribution networks, an attacker can deduce the network topology by accessing the nodal voltage measurements and applying signal processing/machine learning (ML) techniques such as deep neural network [43], linear discriminant analysis and regularized diagonal quadratic discriminant analysis [41], generative adversarial networks [44], and recursive grouping algorithm s[45]. As distribution grids are sparsely monitored, attackers can also use measurements from advanced metering infrastructure (AMI) to derive the network topology [46], [47]. Similar signal processing/ML-based approaches can also be used to deduce the topology of transmission networks [42], [48].
@@
Existing literature often formalizes the attack model of LAAs by co-optimizing it with the distribution system operations as given in and adopting a common cybersecurity practice, which is to conservatively assume a strong (omniscient) attacker and a weak (naive) defender (e.g., DSO) [49, 50, 51]. Accordingly, the omniscient attacker has full knowledge of the system, while the naive defender has no detection or defense mechanisms to respond. These assumptions lead to assessing the worst-case impact of the attack. The omniscient attacker can pursue instant and lasting effects on the power system and DERs, given oracular knowledge of the system conditions. The attacker objective can encompass: (i) reducing efficiency of DSO operations by reducing the energy and cost performance of the system, (ii) equipment damages and failures by incurring additional wear-and-tear due to violating operation limits on equipment and intelligently triggering relay protection to cause load shedding and blackouts, (iii) stressing the system during extreme events (e.g., natural disasters or peak conditions) to increase the likelihood of equipment failures.
@@
The model considered in the works above analyzes the worst-case impact of cyberattacks, i.e., under the assumption of the omniscient attacker and naive defender. In practice, it is unrealistic that an attacker can monitor every aspect of DER or power system operation. Therefore, it is realistic to account for the attacker’s limited knowledge of actions of uncompromised DERs and IoT-loads, and the DSO, which prevents the attacker from precisely estimating attack impacts on the grid and from incurring the theoretical maximum damage. The attacker’s limited knowledge (constraints) can be considered in various manners, including:
@@
(i) Statistical modeling of uncompromised DERs/IoTs: using open-source data sets, e.g., Pecan Street [52] or Tracebase data repositories [53], the attacker can construct a statistical model to describe the behavior of uncompromised devices. These repositories have sufficient data to estimate the mean and standard deviation, which can be used to parameterize the uncompromised devices via the Gaussian distribution, or more complex statistical models (e.g., to account for temporal and spatial correlations [54]). Under such statistical models, the threat model of LAAs can be reformulated into different instances of computationally tractable robust optimization models, which can be used to formalize and compare different attack objectives under various statistical assumptions, i.e., quantify realistic rather than the worst-case attack effects.
@@
(ii) Adversarial Modeling of the Response: Building on statistical models, one can consider an adversarial behavior against the DSO, which renders a game-theoretic formulation for DSO-attacker interactions. This adversarial modeling can provide a realistic rather than naive grid response and can be implemented in a static or dynamic manner, where the attacker fully or partially observes the system response and modifies the attack accordingly [55]. This can determine the existence, uniqueness, and properties of the equilibrium for single/multi-stage (sequential) games using computationally scalable algorithmic game theory [56, 57].
@@
(iii) Learning for Partial System Observability: The statistical and adversarial models described above, can be supported by considering the limited observability of the system by the attacker (e.g., it can only partially observe power flows, voltages, and other system parameters). In this case, the attacker’s problem can be formulated via reinforcement learning (RL), where the observed nodal parameters are the environmental states and the attack damage is the reward. Using such RL frameworks can lead to continuous learning, instead of the omniscient attacker, in a multi-stage adversarial game.
II-B Attack Types
TABLE II: Classification of LAAs.
Criterion Category Description Works
Attack Scope Localized Attacks are designed to specifically target and disrupt a particular region, bus, or feeder within the power grid. [36, 58, 59, 60, 61, 62, 63]
Multi-area Attacks are designed to compromise and manipulate loads across multiple locations. [64, 65, 66, 67, 68, 69, 70, 71, 72, 27, 3, 73, 74, 75, 76]
Feedback Type Static LAAs Attackers alter loads once or at fixed intervals without real-time monitoring of grid response. [36, 76, 62, 72, 27, 3, 64, 65, 61, 66, 63, 67, 68, 69]
Dynamic LAAs Attackers continuously adjust loads in response to system conditions that can be measured [60, 68, 69, 70, 71, 67]
Control Mechanism Informed-Knowledge Attackers possess detailed grid parameters, including load profiles, grid topologies, or system responses. [36, 76, 62, 72, 27, 3, 64, 65, 61, 66, 63, 67, 68, 69, 70, 71, 59, 60, 73, 74, 75, 58]
Limited-Knowledge Attackers have partial or estimated knowledge of the power system, (e.g., frequency measurements). [40]
Targeted System Transmission Grid Attacks aim to destabilize grid frequency or induce cascading failures in the transmission grid. [27, 3, 64, 60, 70, 71, 65, 61, 66, 68, 63, 69, 67, 36, 59, 77]
Distribution Grid Attacks cause localized voltage instability or disrupt demand response programs in the distribution grid. [76, 62, 72]
Energy Markets Attacks manipulate the energy market by artificially inflating or deflating demand for financial gain. [73, 74, 75, 58]
Table II categorizes LAAs based on four key criteria: feedback type, attack scope, control mechanism, and targeted system. This classification provides insight into the diverse strategies used by adversaries and their impact on grid stability and energy markets. The feedback type distinguishes between static LAAs, e.g., [64, 66, 67, 63, 65], where load modifications occur at fixed intervals without real-time monitoring, and dynamic LAAs, e.g., [70, 71, 69], where attackers adjust their actions based on system response, making detection more challenging. The attack scope differentiates between localized attacks, which affect specific buses or feeders, and multi-area attacks, which manipulate loads across multiple regions, potentially causing widespread grid instability. The control mechanism is classified into informed-knowledge attacks, where adversaries have detailed grid parameters, and limited-knowledge attacks, where attackers rely on partial system insights, such as frequency measurements [40]. Finally, the targeted system classification highlights the different objectives of LAAs, including destabilizing the transmission grid by inducing frequency deviations and cascading failures, e.g., [70, 71, 69], disrupting the distribution grid through voltage instability and demand response manipulation [76, 62, 72], and exploiting energy markets by artificially inflating or deflating demand for financial gain [73, 74, 75, 58]. This classification framework underscores the multifaceted nature of LAAs, providing a deeper understanding of their potential impact on grid operations and market dynamics.